OpenStudy (bossimbacon):

What is the slope of the line passing through the points (2, 5) and (0, -4) ___/___

OpenStudy (bossimbacon):

@amistre64

OpenStudy (amistre64):

one option is to play with the slope formula

OpenStudy (lyrae):

Equation of a line: y = mx + b m = slope \[m = \frac{ y_2 - y_1 }{ x_2 - x_1 }\]

OpenStudy (bossimbacon):

i don't understand :/

OpenStudy (amistre64):

how do you define a slope?

OpenStudy (bossimbacon):

i know the formula for slope but i don't really know how to do it :/

OpenStudy (amistre64):

ok, you know the formula ... then we simply have to label the parts

OpenStudy (amistre64):

(2 , 5) and (0, -4) x1 y1 x2 y2 put them in the formula, the rest is adding

OpenStudy (bossimbacon):

so if (mx) would be 2,5? and (0,-4) is b? or something like that

OpenStudy (bossimbacon):

cause the formula is mx+b

OpenStudy (amistre64):

the slope formula is not mx + b

OpenStudy (amistre64):

y1-y2 slope = -------- x1-x2 as previously pointed out

OpenStudy (amistre64):

given 2 points, and label them (2 , 5) and (0, -4) x1 y1 x2 y2 fill in the slope formula

OpenStudy (bossimbacon):

x= 2,5 0,-4?